# Scaling Text-Rich Image Understanding via Code-Guided Synthetic Multimodal Data Generation

### v1

@@ 2 Related Work
Text-rich Images Understanding. Chart understanding and text-rich image understanding continue to challenge state-of-the-art models as naturally occurring vision-language data that can support training for understanding text-rich images is still scarce (Kahou et al., 2017; Kafle et al., 2018; Xu et al., 2023; Mukhopadhyay et al., 2024). In addition to charts and plots, a number of datasets address other kinds of text-rich images such as documents, infographics, diagrams and figures, and screenshots (Siegel et al., 2016; Mathew et al., 2021, 2022; Baechler et al., 2024; Roberts et al., 2024) have been made available. Many of these benchmarks are limited in size and scope, diversity of visualization types, and question types, making them suitable for evaluation but not for training data that could lead to generalized performance.
Figure 2: The overview of our Code Guided Synthetic data generation system (CoSyn), which has 20 generation pipelines based on 11 render tools. Given a user query, e.g., “book cover,” CoSyn selects the appropriate pipelines and starts with generating diverse topics conditioned on personas, then synthesizes detailed data for code generation. The code renders the image and is also fed as context for an LLM to construct instruction-tuning data.
Synthetic Data for VLM. Generating synthetic images with annotations grounded in known source representations has been widely used in domains with limited vision-language data (Johnson-Roberson et al., 2017; Johnson et al., 2017; Cascante-Bonilla et al., 2022; Zhang et al., 2024). This approach has been applied to chart and plot VQA typically using a limited small set of chart types and by instantiating handcrafted question templates (Kahou et al., 2017; Kafle et al., 2018; Methani et al., 2020; Singh and Shekhar, 2020). Following this, Li and Tajbakhsh (2023) and Carbune et al. (2024a) explore using text-only LLMs to generate annotations from tables or text descriptions associated with charts to train VLMs. Other recent approaches, similar to our procedure, explore generating data and code to render synthetic charts Han et al. (2023); Shinoda et al. (2024); Xia et al. (2024). Molmo Deitke et al. (2024) releases a synthetic text-rich image dataset, PixMo-docs, but smaller in scale and diversity than ours. These works generate synthetic data that is still highly limited in terms of the diversity of topics, figure types, and rendering pipelines, which is important for generalizing to out-of-distribution tasks. In our work, we expand the scope beyond charts to encompass a wider range of diverse text-rich images.
@@ 4 CoSyn System
In the following, we provide detailed explanations of the rendering tools supported by CoSyn, the tailored generation pipelines based on these tools, our persona-driven approach to diversify content and styles, and the large-scale dataset of 400K synthetic images generated by CoSyn.
@@
Pipelines. We design 20 pipelines based on 11 rendering tools. [^1] Some tools are used in multiple pipelines, e.g., HTML is used for generating documents, tables, and charts. Each pipeline follows the same procedure: (1) Topic generation to define the theme of this synthetic example, (2) Data generation to populate the detailed contents, (3) Code generation to create executable code that renders the image, and (4) Instruction generation conditioned on code to produce instructions, including questions, answers and explanations for chain-of-thought reasoning. Each stage is controlled by a prompt customized for image category and rendering tool. Figure 8 shows all prompts of the HTML Document pipeline.
@@ 6 Results
Figure 4: Ablation on training data selection. Aux, Syn, and Eval stand for auxiliary, synthetic, and evaluation datasets, respectively. We report the average score on eight benchmarks. The detailed performance breakdown on each benchmark is in Table 7.
@@ 6.2 Analysis
Synthetic data boosts the performance. Table 4 presents an ablation study on the choices of supervised fine-tuning data. In the zero-shot settings, when the model is trained on auxiliary datasets (over 1M training images not directly from the evaluation tasks), it fails to generalize effectively to the evaluation tasks, with a substantial performance gap of 14.1% below GPT-4V. However, using only 400K synthetic samples achieves a performance comparable to GPT-4V. Our best zero-shot model surpasses GPT-4V when jointly training synthetic and auxiliary data. Under the supervised settings, training with in-domain data alone yields strong performance. However, adding 1M auxiliary samples provides a modest improvement of 1.4%, while incorporating synthetic data results in a more significant 3.6% boost. These results demonstrate the effectiveness of synthetic data in enhancing VLMs’ performance on text-rich tasks.
@@
Synthetic Data for Chain-of-Thought Reasoning. Existing text-rich datasets, such as ChartQA Masry et al. (2022), are typically annotated with short answers. However, questions like “Compute the mean of the data in the plot” require step-by-step mathematical reasoning to arrive at the correct answer. Models trained only with short-answer supervision may fail to learn proper plot comprehension, but instead overfitting to annotation biases in these datasets. On the contrary, our CoSyn-400K includes explanation text alongside the short answer. Each instruction-tuning example consists of a (question, explanation, short answer) triplet, enabling models to learn chain-of-thought (CoT) reasoning. During fine-tuning, we design two prompt templates for our synthetic data:
@@
Those prompts allow VLMs to switch between the two answering styles and perform CoT reasoning when necessary. Figure 6 shows that incorporating CoT reasoning improves performance on ChartQA, TableVQA, and NutritionQA, as these datasets contain examples requiring multi-hop reasoning. However, we observe that adding CoT reasoning reduces performance on DocVQA and InfoVQA. We find this decline is caused by answer biases in these benchmarks. Specifically, the ground-truth answers favor short responses, often penalizing more detailed and verbal responses. For instance, in DocVQA, the ground-truth for an example is “T-Th”, whereas the model responds with “Tuesday to Thursday”. Although the response is correct, the strict string-matching metric assigns it a zero score. This highlights key limitations of current multimodal benchmarks, including answering biases and rigid evaluation metrics that fail to capture the full extent of a model’s capabilities.
@@
Similarly, without synthetic data, our model shows a noticeable 21.8% gap between the two question types. However, incorporating synthetic data during training reduces this gap to 14.2%, improving the model’s ability to answer human-asked questions. This suggests that synthetic data can mitigate overfitting on benchmarks and enhance VLMs’ usability in real-world applications.
@@ 6.3 Synthetic Pointing Data
Pointing enables vision-language models to answer questions by providing specific points on images. This functionality allows models to ground their responses in visual content and interact with environments, which is crucial for developing digital agents. We find that we can synthesize pointing data using our code-guided generation system.
@@
Results. Table 3 compares the click accuracy of our models with previous methods. Using 65K synthetic pointing samples, our model achieves performance comparable to the one trained on 155K human-annotated samples. When combining synthetic and human data, our model achieves state-of-the-art performance on ScreenSpot, surpassing the recent UGround Gou et al. (2024), which was trained on 1.3M screenshots. These results demonstrate that synthetic pointing data is a data-efficient approach for improving VLM performance on agentic tasks involving click prediction.
@@ A.4 Academic Datasets
A.5 Training Details
Hyper Parameters. We set the maximum sequence length for training is 2304 tokens. We use the same learning rate of 1e-6 for the MLP connector, LLM, and visual encoder, with batch size 32. The best-performing model is trained for 60K steps with 200 warm-up steps and a cosine scheduler with an end factor of 0.1. All experiments are run on a single TPU v3-128.
Appendix B Additional Analysis
Dataset Image Diversity Text Diversity
FigureQA 0.268 0.567
DVQA 0.307 0.752
PlotQA 0.420 0.743
ChartQA 0.340 0.742
Ours (Charts) 0.596 0.823
n. of Tools Diversity ChartQA
Average Machine Human
Single 0.572 73.9 66.5 81.5
Multiple 0.607 75.2 68.6 82.0
@@ Appendix B Additional Analysis
LLM for Data Generation ChartQA
Average Machine Human
GPT-4o 72.4 65.8 78.9
Claude-3.5-sonnet 77.2 71.0 83.8
@@ Appendix C Qualitative Examples
Use of AI Assistants. We use AI to fix some typos and grammar. Authors write all contents.

### v2

@@ 2 Related Work
Text-rich Images Understanding. Chart understanding and text-rich image understanding continue to challenge state-of-the-art models as naturally occurring vision-language data that can support training for understanding text-rich images is still scarce (Kahou et al., 2017; Kafle et al., 2018; Xu et al., 2023; Mukhopadhyay et al., 2024). In addition to charts and plots, a number of datasets address other kinds of text-rich images such as documents, infographics, diagrams, and figures, and screenshots (Siegel et al., 2016; Mathew et al., 2021, 2022; Baechler et al., 2024; Roberts et al., 2024) have been made available. Many of these benchmarks are limited in size and scope, diversity of visualization types, and question types, making them suitable for evaluation but not for training data that could lead to generalized performance.
Figure 2: The overview of our Code Guided Synthetic data generation system (CoSyn), which has 20 generation pipelines based on 11 render tools. Given a user query, e.g., “book cover,” CoSyn selects the appropriate pipelines and starts by generating diverse topics conditioned on personas, then synthesizes detailed data for code generation. The code renders the image and is also fed as context for an LLM to construct instruction-tuning data.
Synthetic Data for VLM. Generating synthetic images with annotations grounded in known source representations has been widely used in domains with limited vision-language data (Johnson-Roberson et al., 2017; Johnson et al., 2017; Cascante-Bonilla et al., 2022; Zhang et al., 2024). This approach has been applied to chart and plot VQA, typically using a limited, small set of chart types and by instantiating handcrafted question templates (Kahou et al., 2017; Kafle et al., 2018; Methani et al., 2020; Singh and Shekhar, 2020). Following this, Li and Tajbakhsh (2023) and Carbune et al. (2024a) explore using text-only LLMs to generate annotations from tables or text descriptions associated with charts to train VLMs. Other recent approaches, similar to our procedure, explore generating data and code to render synthetic charts Han et al. (2023); He et al. (2024); Shinoda et al. (2024); Xia et al. (2024). Molmo Deitke et al. (2024) releases a synthetic text-rich image dataset, PixMo-docs, but smaller in scale and diversity than ours. These works generate synthetic data that is still highly limited in terms of the diversity of topics, figure types, and rendering pipelines, which is important for generalizing to out-of-distribution tasks. We expand the scope beyond charts to encompass a wider range of diverse text-rich images.
@@ 4 CoSyn System
In the following, we provide detailed explanations of the rendering tools supported by CoSyn, the tailored generation pipelines based on these tools, our persona-driven approach to diversifying content and styles, and the large-scale dataset of 400K synthetic images generated by CoSyn.
@@
Pipelines. We design 20 pipelines based on 11 rendering tools. [^1] Some tools are used in multiple pipelines, e.g., HTML is used for generating documents, tables, and charts. Each pipeline follows the same procedure: (1) Topic generation to define the theme of this synthetic example, (2) Data generation to populate the detailed contents, (3) Code generation to create executable code that renders the image, and (4) Instruction generation conditioned on code to produce instructions, including questions, answers and explanations for chain-of-thought reasoning. Each stage is controlled by a prompt customized for the image category and rendering tool. Figure 9 shows prompts of the HTML Document pipeline.
@@ 6 Results
Figure 4: Ablation on training data selection. Aux, Syn, and Eval stand for auxiliary, synthetic, and evaluation datasets, respectively. We report the average score on seven benchmarks. The detailed performance breakdown on each benchmark is in Table 8.
@@ 6.2 Analysis
Synthetic data boosts the performance. Figure 4 presents an ablation study on the choices of supervised fine-tuning data. In the zero-shot settings, when the model is trained on auxiliary datasets (over 1M training images not directly from the evaluation tasks), it fails to generalize effectively to the evaluation tasks, with a substantial performance gap of 14.1% below GPT-4V. However, using only 400K synthetic samples achieves a performance comparable to GPT-4V. Our best zero-shot model surpasses GPT-4V when jointly training synthetic and auxiliary data. Under the supervised settings, training with in-domain data alone yields strong performance. However, adding 1M auxiliary samples provides a modest improvement of 1.4%, while incorporating synthetic data results in a more significant 3.6% boost. These results demonstrate the effectiveness of synthetic data in enhancing VLMs’ performance on text-rich tasks.
@@
Synthetic Data for Chain-of-Thought Reasoning. Existing text-rich datasets, such as ChartQA Masry et al. (2022), are typically annotated with short answers. However, questions like “Compute the mean of the data in the plot” require step-by-step mathematical reasoning to arrive at the correct answer. Models trained only with short-answer supervision may fail to learn proper plot comprehension, but instead overfit to annotation biases in these datasets. On the contrary, our CoSyn-400K includes explanation text alongside the short answer. Each instruction-tuning example consists of a (question, explanation, short answer) triplet, enabling models to learn chain-of-thought (CoT) reasoning. During fine-tuning, we design two prompt templates for our synthetic data:
@@
Those prompts allow VLMs to switch between the two answering styles and perform CoT reasoning when necessary. Figure 6 shows that incorporating CoT reasoning improves performance on ChartQA, TableVQA, and NutritionQA, as these datasets contain examples requiring multi-hop reasoning. However, we observe that adding CoT reasoning reduces performance on DocVQA and InfoVQA. We find that answer biases in these benchmarks cause this decline. Specifically, the ground-truth answers favor short responses, often penalizing more detailed and verbal responses. For instance, in DocVQA, the ground-truth for an example is “T-Th”, whereas the model responds with “Tuesday to Thursday”. Although the response is correct, the strict string-matching metric assigns it a zero score. This highlights key limitations of current multimodal benchmarks, including answering biases and rigid evaluation metrics that fail to capture the full extent of a model’s capabilities.
@@
Similarly, without synthetic data, our model shows a noticeable 21.8% gap between the two question types. However, incorporating synthetic data during training reduces this gap to 14.2%. This suggests that synthetic data can mitigate overfitting on benchmarks and enhance VLMs’ usability in real-world human-asked questions.
Dataset Image Diversity Text Diversity
FigureQA 0.268 0.567
DVQA 0.307 0.752
PlotQA 0.420 0.743
ChartQA 0.340 0.742
Ours (Charts) 0.596 0.823
n. of Tools Diversity ChartQA
Average Machine Human
Single 0.572 73.9 66.5 81.5
Multiple 0.607 75.2 68.6 82.0
LLM for Data Generation ChartQA
Average Machine Human
GPT-4o 72.4 65.8 78.9
Claude-3.5-sonnet 77.2 71.0 83.8
@@ 6.3 Synthetic Pointing Data
Pointing enables VLMs to answer questions by providing specific points on images. This allows models to ground their responses in visual content and interact with environments, which is crucial for developing digital agents. We find that our code-guided system can synthesize pointing data.
@@
Results. Table 6 compares the click accuracy of our models with previous methods. Using 65K synthetic pointing samples, our model achieves performance comparable to that trained on 155K human-annotated samples. When combining synthetic and human data, our model achieves state-of-the-art performance on ScreenSpot, surpassing the recent UGround Gou et al. (2024), which was trained on 1.3M screenshots. These results demonstrate that synthetic pointing data is a data-efficient approach for improving VLM performance on agentic tasks involving click prediction.
@@ A.4 Academic Datasets
A.5 Training Details
Hyper Parameters. We set the maximum sequence length for training to 2304 tokens. We use the same learning rate of 1e-6 for the MLP connector, LLM, and visual encoder, with batch size 32. The best-performing model is trained for 60K steps with 200 warm-up steps and a cosine scheduler with an end factor of 0.1. All experiments are run on a single TPU v3-128.
Appendix B Additional Analysis